OpenStudy (anonymous):

Write the equation of the circle with center (-2, -5) and containing the point (-10, -20). Use the ^ key for the exponents. Write your answer as the example: (x+7)^2+(y-8)^2=100

OpenStudy (anonymous):

the equation of the circle is \[(x-h)^2+(y-k)^2=r^2\] where: (h,k) Center and r the radio replace the center and you will have: \[(x+2)^2+(y+5)^2=r^2\] Now, you need find the radio so you have to use the points given. Replace the points and you have the radio

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

how to find their distance ? well \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-2\quad ,&-5)\quad &(-10\quad ,&-20) \end{array}\qquad d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (anonymous):

ok... i'm still confused no offence to any of you.

OpenStudy (amigatour):

(x−h)^2+(y−k)^2=r2 (x−4)^2+(y+4)^2=180 How did I get the radius? Well, I used the distance formula: check Jdoe001's work . The actual distance is \[\sqrt{180}\] or ~13.4 13.4^2=? It's about 180.